# Point-Level Topological Representation Learning on Point Clouds

### v2

@@ Abstract
Figure 1: Computing Topological Point Features (topf). Input. A point cloud $X$ in $n$-dimensional space. Step 1. To extract global topological information, the persistent homology is computed on an $\alpha$/vr-filtration. The most significant topological features $\mathcal{F}$ across all specified dimensions are selected. Step 2. $k$-homology generators associated to all features $f_{i,k}\in\mathcal{F}$ are computed. For every feature, a simplicial complex is built at a step of the filtration where $f_{i,k}$ is alive. Step 3. The homology generators are projected to the harmonic space. Step 4. The vectors are normalised to obtain vectors $\smash{\mathbf{e_{k}^{i}}}$ indexed over the $k$-simplices. For every point $x$ and feature $\smash{f\in\mathcal{F}}$, we compute the mean of the entries of $\mathbf{e_{k}^{i}}$ corresponding to simplices containing $x$. The output is a $|X|\times|\mathcal{F}|$ matrix which can be used for downstream ml tasks. Optional. We weigh the simplicial complexes resulting in a topologically more faithful harmonic representative in Step 3.
@@ 2 Main Ideas of Topf
$\smash{L_{n}\coloneq\mathcal{B}_{n-1}^{\top}\mathcal{B}_{n-1}+\mathcal{B}_{n}\mathcal{B}_{n}^{\top}.}$
@@ 3 How to Compute Topological Point Features
We now need to select the relevant persistent homology classes which carry the most important global information. The persistent homology $P_{k}$ module in dimension $k$ is given to us as a list of pairs of birth and death times $(b_{i}^{k},d_{i}^{k})$. We can assume these pairs are ordered in non-increasing order of the durations $l_{i}^{k}=d_{i}^{k}-b_{i}^{k}$. This list is typically very long and consists to a large part of noisy homological features which vanish right after they appear. In contrast, we are interested in connected components, loops, cavities, etc. that persist over a long time, indicating that they are important for the shape of the point cloud. Distinguishing between the relevant and the irrelevant features is in general difficult and may depend on additional insights on the domain of application. In order to provide a heuristic which does not depend on any a-priori assumptions on the number of relevant features we pick the smallest quotient $q_{i}^{k}\coloneq l_{i+1}^{k}/l_{i}^{k}>0$ as the point of cut-off $\smash{N_{k}\coloneq\operatorname*{arg\,min}_{i}q_{i}^{k}}$. The only underlying assumption of this approach is that the band of ‘relevant’ features is separated from the ‘noisy’ homological features by a drop in persistence. If this assumption is violated, the only possible way to do meaningful feature selection depends on application-specific domain knowledge. We found that our proposed heuristics work well across a large scale of applications. See Figure 2 left and centre for an illustration and Appendix G for more technical details and ways to improve and adapt the feature selection module of topf. We call the chosen $k$-homology classes with $k$-homology generators $f^{i}_{k}$.
@@
In this step, we need to relate the global topology extracted in the previous step to the simplices which we will use to compute the local topological point features. Every selected feature $f_{k}^{i}$ of the previous step comes with a birth time $b_{i,k}$ and a death time $d_{i,k}$. This means that the homology class $f_{k}^{i}$ is present in every sc of the filtration between step $\varepsilon=b_{i,k}$ and $\varepsilon=d_{i,k}$ and we could choose any of the scs for the next step. Picking a small $\varepsilon$ will lead to fewer simplices in the sc and thus to a very localised harmonic representative. Picking a large $\varepsilon$ will lead to many simplices in the sc and thus to a very smooth and ‘blurry’ harmonic representative with large support. Finding a middle ground between these regimes returns optimal results. Given interpolation hyperparameter $\gamma\in(0,1)$, we will thus consider the simplicial complex $\mathcal{S}^{t_{i,k}}(X)$ at step $\smash{t_{i,k}\coloneq b_{i,k}^{1-\gamma}d_{i,k}^{\gamma}}$ for $k>0$ and at step $t_{i,k}\coloneq\gamma d_{i,k}$ for $k=0$ of the simplicial filtration. At this point, the homology class $f_{k}^{i}$ is still alive. We then consider the real vector space $\smash{\mathbb{R}[\mathcal{S}_{k}^{t_{i,k}}(X)]}$ with formal basis consisting of the $k$-simplices of the sc $\mathcal{S}^{t_{i,k}}$. From the persistent homology computation of the first step, we also obtain a generator of the feature $f_{k}^{i}$, consisting of a list $\Sigma_{k}^{i}$ of simplices $\smash{\hat{\sigma}_{j}\in\mathcal{S}^{b_{i,k}}_{k}}$ and coefficients $c_{j}\in\mathbb{Z}/3\mathbb{Z}$. We need to turn this formal sum of simplices with $\mathbb{Z}/3\mathbb{Z}$-coefficients into a vector in the real vector space $\smash{\mathbb{R}[\mathcal{S}_{k}^{t_{i,k}}(X)]}$: Let $\iota\colon\mathbb{Z}/3\mathbb{Z}$ be the map induced by the canonical inclusion of $\{-1,0,1\}\hookrightarrow\mathbb{R}$. We can now define an indicator vector $\smash{e_{k}^{i}\in\mathbb{R}[\mathcal{S}_{k}^{t_{i,k}}(X)]}$ associated to the feature $f_{k}^{i}$ (Cf. DeSilva2009persistent).
@@
Empirically, $e^{i}_{k}$ is a homology generator for all real coefficients as well, although our construction only guarantees for $\mathcal{B}_{k-1}e^{i}_{k}\equiv 0\mod 3$. We discuss how to fix the rare case where this does not work in Section B.3 and now assume to work with a real homology representative $e^{i}_{k}$. While this homology representative lives in a real vector space, it is not unique, has a small support, and its value can differ largely even between close simplices. All of these problems can be solved by projecting the homology representative to the harmonic subspace $\ker L_{k}$ of $\smash{\mathbb{R}[\mathcal{S}_{k}^{t_{i,k}}(X)]}$. Rather than directly projecting $e_{k}^{i}$ to the harmonic subspace, we make use of the Hodge decomposition theorem (Theorem 2.5) which allows us to solve computationally efficient least square problems:
@@
and get the harmonic representative $\hat{e}_{k}^{i}\coloneqq e_{k}^{i}-e_{k,\text{grad}}^{i}-e_{k,\text{curl}}^{i}$. (Cf. Figure 2 right for a visualisation.) Because homology representatives are gradient-free, $\smash{e_{k,\text{grad}}^{i}=0}$ and we only need to compute $\smash{e_{k,\text{curl}}^{i}}$.
@@
In the previous step, we have computed a set of harmonic representatives of homology classes. Each harmonic representative assigns each simplex in the correspondings scs a value. However, these simplices likely have no real-world meaning and the underlying simplicical complexes differ depending on the birth and death times of the homology classes. Hence in this step, we will collect the features on the point-level after performing some necessary preprocessing. Given a vector $\hat{e}_{k}^{i}$ indexed over simplices and a hyperparameter $\delta$, we now construct $\mathbf{e}_{k}^{i}\colon\mathcal{S}_{k}^{t_{i,k}}(X)\rightarrow[0,1]$ by setting $\mathbf{e}_{k}^{i}\colon\sigma\mapsto\in\{|\hat{e}_{k}^{i}(\sigma)|/(\delta\max_{\sigma^{\prime}\in\mathcal{S}_{k}^{t_{i,k}}(X)}|\hat{e}_{k}^{i}(\sigma^{\prime})|),1\}$ such that $\hat{e}_{k}^{i}$ is normalised to $\smash{[0,1]}$, then the values of $\smash{[0,\delta]}$ are mapped linearly to $\smash{[0,1]}$ and everything above is sent to $1$. We found empirically that a thresholding parameter of $\delta\approx 0.07$ works best across at the range of applications considered below. However, topf is not sensitive to small changes to $\delta$ because entries of $\hat{e}_{k}^{i}$ are concentrated around 00 (cf. Section E.1).
@@ 4 Experiments
Figure 3: topf on 3d real-world and synthetic point clouds. For every point, we highlight the largest corresponding topological feature, where colour stands for the different features and saturation for the value of the feature. Formulate nicer?). (b): Atoms of nalcn vhannelosome (Kschonsak:2022) display three distinct loops. (c): Points sampled in the state space of a Lorentz attractor. The two features correspond to the two lobes of the attractor. (d): Point cloud spaceship of our newly introduced topological clustering benchmark suite (See Appendix D). (e): Latent space of a vae trained on image patches showing topological structure (See Figure 12 for details).
@@
Table 1: Quantitative performance comparison of clustering with topf and other features/clustering algorithms. Four 2d and three 3d data sets of the topological clustering benchmark suite (Appendix D, cf. Figure 8 for ground truth and Figure 9 for labels by topf). We ran each algorithm $20$ times and list the mean adjusted rand index (ari) with standard deviation $\sigma$ and mean running time. We omit $\sigma$ for algorithms with $\sigma=0$. Spectral clustering on topf vectors consistently outperforms or almost matches the other algorithms while having significantly better run time than the second best performing algorithm tpcc. Spectral Clustering (sc), dbscan, and Agglomerative Clustering (AgC) (applied on the points) are standard clustering algorithms, tomato is a topological clustering algorithm (Chazal2013), Geo clusters using $12$-dimensional point geometric features extracted by pgeof, whereas node2vec (grover2016node2vec) produces node embeddings on a $k$-nearest neighbour graph. We pretrained PointNet (pn) (qi2017pointnet) on the ShapeNetPart point cloud segmentation data set (chang2015shapenet). ‘Weakly Supervised 3D Local Descriptor Learning for Point Cloud Registration’ (wsd (Li2022WSDesc)) is pretrained on 3DMatch data and produces $32$-dimensional feature vectors.We highlight all ari scores within ±$\pm 0.05$ of the best ari score.
topf tpcc sc dbscan tomato Geo node2vec pn wsd
4spheres ari 0.81 0.52$\pm$0.17 0.37 0.00 0.32 0.20 0.00$\pm$0.00 0.30 0.13
time (s) 13.9 23.3 0.2 <0.1 <0.1 0.2 48.4 <0.1 <0.1
Ellipses ari 0.95 0.47$\pm$0.04 0.25 0.19 0.29 0.81 0.02$\pm$0.00 0.50 0.35
time (s) 13.5 14.4 0.1 <0.1 <0.1 0.1 11.2 <0.1 <0.1
4Circles+Grid ari 0.70 0.39$\pm$0.04 0.90 0.92 0.82 0.41 0.01$\pm$0.00 0.55 0.06
time (s) 13.7 28.5 0.5 <0.1 <0.1 0.3 63.8 <0.1 <0.1
Halved Circle ari 0.71 0.18$\pm$0.12 0.24 0.00 0.16 0.08 0.00$\pm$0.01 0.36 0.00
time (s) 13.6 14.3 0.1 <0.1 <0.1 0.1 18.2 <0.1 <0.1
2Spheres2Circle ari 0.94 0.97$\pm$0.01 0.70 0.00 0.87 0.12 0.00$\pm$0.00 0.55 0.95
time (s) 20.7 1662.2 1.6 <0.1 <0.1 0.9 348.6 <0.1 <0.1
SphereinCircle ari 0.97 0.98$\pm$<0.1 0.34 0.00 0.06 0.69 0.13$\pm$0.03 0.39 0.46
time (s) 14.5 8.0 <0.1 <0.1 <0.1 0.08 20.1 <0.1 <0.1
Spaceship ari 0.92 0.56$\pm$0.03 0.28 0.26 0.30 0.87 0.07$\pm$0.00 0.41 0.76
time (s) 15.5 341.8 16.7 <0.1 <0.1 0.2 49.8 <0.1 <0.1
mean ari 0.86 0.58 0.44 0.16 0.40 0.45 0.03 0.44 0.39
time (s) 15.1 298.9 0.4 <0.1 <0.1 0.3 80.0 <0.1 <0.1
@@
In Figure 3, we show qualitatively that topf constructs meaningful topological features on data sets from Biology and Physics, and synthetic data, corresponding to for example rings and pockets in proteins or trajectories around different attractors in dynamical systems, (for individual heatmaps see Figure 14). In Figure 4, we show that topf works for a high-dimensional conformation space.
@@ B.1 More details on vr and $\alpha$-filtrations
Vietoris–Rips complexes are easy to define, approximate the topological properties of a point cloud across all scales and computationally easy to implement. However for moderately large $r$, the associated vr complex contains a large number of simplices — up to $\binom{|X|}{n+1}$ $n$-simplices for large enough $r$ — leading to poor computational performance for any downstream task on some large point clouds. One way to see this is the following: After adding the first edge that connects two components or the final simplex that fills a hole in the simplicial complex the vr complex keeps adding more and more simplices in the same area that keep the topology unchanged. One way to mitigate this problem is to pre-compute a set of simplices that are able to express the entire topology of the point cloud. For a point cloud $X\subset\mathbb{R}^{n}$, the $\alpha$-filtration consists of the intersection of the simplicial complexes of the vr filtration on $X$ with the (higher-dimensional) Delaunay triangulation of $X$ in $\mathbb{R}$. Due to algorithmic reasons, the filtration value of a simplex is then related to the radius of the circumscribed sphere instead of the maximum pair-wise distance of vertices, where the filtration value $\alpha^{X}(\sigma)$ of a simplex $\sigma$ is given by the minimum $\varepsilon$ such that $\sigma$ is contained in $\alpha_{\varepsilon}(X)$, i.e. $\alpha^{X}(\sigma)\coloneqq\inf\{\varepsilon\in\mathbb{R}:\sigma\in\alpha_{\varepsilon}(V)\}$. This reduces the number of required simplices across all dimensions to $\smash{O(|X|^{\lceil n/2\rceil})}$. However, the Delaunay triangulation becomes computationally infeasible for larger $n$.

### v3

@@ Abstract
Figure 1: Computing Topological Point Features (topf). Input. A point cloud $X$ in $n$-dimensional space. Step 1. To extract global topological information, the persistent homology is computed on an $\alpha$/vr-filtration. The most significant topological features $\mathcal{F}$ across all specified dimensions are selected. Step 2. $k$-homology generators associated to all features $f_{i,k}\in\mathcal{F}$ are computed. For every feature, a simplicial complex is built at a step of the filtration where $f_{i,k}$ is alive. Step 3. The homology generators are projected to the harmonic space. Step 4. The vectors are normalised to obtain vectors $\mathbf{e_{k}^{i}}$ indexed over the $k$-simplices. For every point $x$ and feature $f\in\mathcal{F}$, we compute the mean of the entries of $\mathbf{e_{k}^{i}}$ corresponding to simplices containing $x$. The output is a $|X|\times|\mathcal{F}|$ matrix which can be used for downstream ml tasks. Optional. We weigh the simplicial complexes resulting in a topologically more faithful harmonic representative in Step 3.
@@ 2 Main Ideas of topf
$L_{n}\coloneq\mathcal{B}_{n-1}^{\top}\mathcal{B}_{n-1}+\mathcal{B}_{n}\mathcal{B}_{n}^{\top}.$
@@ 3 How to Compute Topological Point Features
We now need to select the relevant persistent homology classes which carry the most important global information. The persistent homology $P_{k}$ module in dimension $k$ is given to us as a list of pairs of birth and death times $(b_{i}^{k},d_{i}^{k})$. We can assume these pairs are ordered in non-increasing order of the durations $l_{i}^{k}=d_{i}^{k}-b_{i}^{k}$. This list is typically very long and consists to a large part of noisy homological features which vanish right after they appear. In contrast, we are interested in connected components, loops, cavities, etc. that persist over a long time, indicating that they are important for the shape of the point cloud. Distinguishing between the relevant and the irrelevant features is in general difficult and may depend on additional insights on the domain of application. In order to provide a heuristic which does not depend on any a-priori assumptions on the number of relevant features we pick the smallest quotient $q_{i}^{k}\coloneq l_{i+1}^{k}/l_{i}^{k}>0$ as the point of cut-off $N_{k}\coloneq\operatorname*{arg\,min}_{i}q_{i}^{k}$. The only underlying assumption of this approach is that the band of ‘relevant’ features is separated from the ‘noisy’ homological features by a drop in persistence. If this assumption is violated, the only possible way to do meaningful feature selection depends on application-specific domain knowledge. We found that our proposed heuristics work well across a large scale of applications. See Figure 2 left and centre for an illustration and Appendix G for more technical details and ways to improve and adapt the feature selection module of topf. We call the chosen $k$-homology classes with $k$-homology generators $f^{i}_{k}$.
@@
In this step, we need to relate the global topology extracted in the previous step to the simplices which we will use to compute the local topological point features. Every selected feature $f_{k}^{i}$ of the previous step comes with a birth time $b_{i,k}$ and a death time $d_{i,k}$. This means that the homology class $f_{k}^{i}$ is present in every sc of the filtration between step $\varepsilon=b_{i,k}$ and $\varepsilon=d_{i,k}$ and we could choose any of the scs for the next step. Picking a small $\varepsilon$ will lead to fewer simplices in the sc and thus to a very localised harmonic representative. Picking a large $\varepsilon$ will lead to many simplices in the sc and thus to a very smooth and ‘blurry’ harmonic representative with large support. Finding a middle ground between these regimes returns optimal results. Given interpolation hyperparameter $\gamma\in(0,1)$, we will thus consider the simplicial complex $\mathcal{S}^{t_{i,k}}(X)$ at step $t_{i,k}\coloneq b_{i,k}^{1-\gamma}d_{i,k}^{\gamma}$ for $k>0$ and at step $t_{i,k}\coloneq\gamma d_{i,k}$ for $k=0$ of the simplicial filtration. At this point, the homology class $f_{k}^{i}$ is still alive. We then consider the real vector space $\mathbb{R}[\mathcal{S}_{k}^{t_{i,k}}(X)]$ with formal basis consisting of the $k$-simplices of the sc $\mathcal{S}^{t_{i,k}}$. From the persistent homology computation of the first step, we also obtain a generator of the feature $f_{k}^{i}$, consisting of a list $\Sigma_{k}^{i}$ of simplices $\hat{\sigma}_{j}\in\mathcal{S}^{b_{i,k}}_{k}$ and coefficients $c_{j}\in\mathbb{Z}/3\mathbb{Z}$. We need to turn this formal sum of simplices with $\mathbb{Z}/3\mathbb{Z}$-coefficients into a vector in the real vector space $\mathbb{R}[\mathcal{S}_{k}^{t_{i,k}}(X)]$: Let $\iota\colon\mathbb{Z}/3\mathbb{Z}$ be the map induced by the canonical inclusion of $\{-1,0,1\}\hookrightarrow\mathbb{R}$. We can now define an indicator vector $e_{k}^{i}\in\mathbb{R}[\mathcal{S}_{k}^{t_{i,k}}(X)]$ associated to the feature $f_{k}^{i}$ (Cf. DeSilva2009persistent).
@@
Empirically, $e^{i}_{k}$ is a homology generator for all real coefficients as well, although our construction only guarantees for $\mathcal{B}_{k-1}e^{i}_{k}\equiv 0\mod 3$. We discuss how to fix the rare case where this does not work in Section B.3 and now assume to work with a real homology representative $e^{i}_{k}$. While this homology representative lives in a real vector space, it is not unique, has a small support, and its value can differ largely even between close simplices. All of these problems can be solved by projecting the homology representative to the harmonic subspace $\ker L_{k}$ of $\mathbb{R}[\mathcal{S}_{k}^{t_{i,k}}(X)]$. Rather than directly projecting $e_{k}^{i}$ to the harmonic subspace, we make use of the Hodge decomposition theorem (Theorem 2.5) which allows us to solve computationally efficient least square problems:
@@
and get the harmonic representative $\hat{e}_{k}^{i}\coloneqq e_{k}^{i}-e_{k,\text{grad}}^{i}-e_{k,\text{curl}}^{i}$. (Cf. Figure 2 right for a visualisation.) Because homology representatives are gradient-free, $e_{k,\text{grad}}^{i}=0$ and we only need to compute $e_{k,\text{curl}}^{i}$.
@@
In the previous step, we have computed a set of harmonic representatives of homology classes. Each harmonic representative assigns each simplex in the correspondings scs a value. However, these simplices likely have no real-world meaning and the underlying simplicical complexes differ depending on the birth and death times of the homology classes. Hence in this step, we will collect the features on the point-level after performing some necessary preprocessing. Given a vector $\hat{e}_{k}^{i}$ indexed over simplices and a hyperparameter $\delta$, we now construct $\mathbf{e}_{k}^{i}\colon\mathcal{S}_{k}^{t_{i,k}}(X)\rightarrow[0,1]$ by setting $\mathbf{e}_{k}^{i}\colon\sigma\mapsto\in\{|\hat{e}_{k}^{i}(\sigma)|/(\delta\max_{\sigma^{\prime}\in\mathcal{S}_{k}^{t_{i,k}}(X)}|\hat{e}_{k}^{i}(\sigma^{\prime})|),1\}$ such that $\hat{e}_{k}^{i}$ is normalised to $[0,1]$, then the values of $[0,\delta]$ are mapped linearly to $[0,1]$ and everything above is sent to $1$. We found empirically that a thresholding parameter of $\delta\approx 0.07$ works best across at the range of applications considered below. However, topf is not sensitive to small changes to $\delta$ because entries of $\hat{e}_{k}^{i}$ are concentrated around 00 (cf. Section E.1).
@@ 4 Experiments
Figure 3: topf on 3d real-world and synthetic point clouds. For every point, we highlight the largest corresponding topological feature, where colour stands for the different features and saturation for the value of the feature. (b): Atoms of nalcn channelosome (Kschonsak:2022) display three distinct loops. (c): Points sampled in the state space of a Lorentz attractor. The two features correspond to the two lobes of the attractor. (d): Point cloud spaceship of our newly introduced topological clustering benchmark suite (See Appendix D). (e): Latent space of a vae trained on image patches showing topological structure (See Figure 12 for details).
@@
Table 1: Quantitative performance comparison of clustering with topf and other features/clustering algorithms. Four 2d and three 3d data sets of the topological clustering benchmark suite (Appendix D, cf. Figure 8 for ground truth and Figure 9 for labels by topf). We ran each algorithm $20$ times and list the mean adjusted rand index (ari) with standard deviation $\sigma$ and mean running time. We omit $\sigma$ for algorithms with $\sigma=0$. Spectral clustering on topf vectors consistently outperforms or almost matches the other algorithms while having significantly better run time than the second best performing algorithm tpcc. Spectral Clustering (sc), dbscan, and Agglomerative Clustering (AgC) (applied on the points) are standard clustering algorithms, tomato is a topological clustering algorithm (Chazal2013), Geo clusters using $12$-dimensional point geometric features extracted by pgeof. We pretrained PointNet (pn) (qi2017pointnet) on the ShapeNetPart point cloud segmentation data set (chang2015shapenet). ‘Weakly Supervised 3D Local Descriptor Learning for Point Cloud Registration’ (wsd (Li2022WSDesc)) is pretrained on 3DMatch data and produces $32$-dimensional feature vectors. dgcnn wasserman2018topological is pretrained on ShapeNetPart segmentation, with the parameters determined by a hyperparameter search. We highlight all ari scores within ±$\pm 0.05$ of the best ari score.
TOPF tpcc sc dbscan tomato Geo pn wsd dgcnn
4spheres ari 0.81 0.52$\pm$0.17 0.37 0.00 0.32 0.20 0.30 0.13 0.29$\pm$0.03
time (s) 13.9 23.3 0.2 <0.1 <0.1 0.2 <0.1 <0.1 <0.1
Ellipses ari 0.95 0.47$\pm$0.04 0.25 0.19 0.29 0.81 0.50 0.35 0.70$\pm$0.00
time (s) 13.5 14.4 0.1 <0.1 <0.1 0.1 <0.1 <0.1 <0.1
4Circles+Grid ari 0.70 0.39$\pm$0.04 0.90 0.92 0.82 0.41 0.55 0.06 0.75$\pm$0
time (s) 13.7 28.5 0.5 <0.1 <0.1 0.3 <0.1 <0.1 <0.1
Halved Circle ari 0.71 0.18$\pm$0.12 0.24 0.00 0.16 0.08 0.36 0.00 0.39$\pm$0.07
time (s) 13.6 14.3 0.1 <0.1 <0.1 0.1 <0.1 <0.1 <0.1
2Spheres2Circle ari 0.94 0.97$\pm$0.01 0.70 0.00 0.87 0.12 0.55 0.95 0.84$\pm$0.18
time (s) 20.7 1662.2 1.6 <0.1 <0.1 0.9 <0.1 <0.1 0.2
SphereinCircle ari 0.97 0.98$\pm$<0.1 0.34 0.00 0.06 0.69 0.39 0.46 0.76$\pm$0.06
time (s) 14.5 8.0 <0.1 <0.1 <0.1 0.08 <0.1 <0.1 <0.1
Spaceship ari 0.92 0.56$\pm$0.03 0.28 0.26 0.30 0.87 0.41 0.76 0.79$\pm$0.00
time (s) 15.5 341.8 16.7 <0.1 <0.1 0.2 <0.1 <0.1 <0.1
mean ari 0.86 0.58 0.44 0.16 0.40 0.45 0.44 0.39 0.64
time (s) 15.1 298.9 0.4 <0.1 <0.1 0.3 <0.1 <0.1 <0.1
@@
In Figure 3, we show qualitatively that topf constructs meaningful topological features on data sets from Biology and Physics, and synthetic data, corresponding to for example rings and pockets in proteins or trajectories around different attractors in dynamical systems, (for individual heatmaps see Figure 14). In Figure 4, we show that topf works for the high-dimensional conformation space of cyclooctane.
@@ B.1 More details on vr and $\alpha$-filtrations
Vietoris–Rips complexes are easy to define, approximate the topological properties of a point cloud across all scales and computationally easy to implement. However for moderately large $r$, the associated vr complex contains a large number of simplices — up to $\binom{|X|}{n+1}$ $n$-simplices for large enough $r$ — leading to poor computational performance for any downstream task on some large point clouds. One way to see this is the following: After adding the first edge that connects two components or the final simplex that fills a hole in the simplicial complex the vr complex keeps adding more and more simplices in the same area that keep the topology unchanged. One way to mitigate this problem is to pre-compute a set of simplices that are able to express the entire topology of the point cloud. For a point cloud $X\subset\mathbb{R}^{n}$, the $\alpha$-filtration consists of the intersection of the simplicial complexes of the vr filtration on $X$ with the (higher-dimensional) Delaunay triangulation of $X$ in $\mathbb{R}$. Due to algorithmic reasons, the filtration value of a simplex is then related to the radius of the circumscribed sphere instead of the maximum pair-wise distance of vertices, where the filtration value $\alpha^{X}(\sigma)$ of a simplex $\sigma$ is given by the minimum $\varepsilon$ such that $\sigma$ is contained in $\alpha_{\varepsilon}(X)$, i.e. $\alpha^{X}(\sigma)\coloneqq\inf\{\varepsilon\in\mathbb{R}:\sigma\in\alpha_{\varepsilon}(V)\}$. This reduces the number of required simplices across all dimensions to $O(|X|^{\lceil n/2\rceil})$. However, the Delaunay triangulation becomes computationally infeasible for larger $n$.